OpenStudy (anonymous):

Solve the equation... 9 over y + 5 - 6 over y-5=6 over y^2-25

jhonyy9 (jhonyy9):

will be 9(y-5)-6(y+5) ------------- = 6/(y2 -25) y2 -25 9y-45-6y-30=6 3y=-69 y= -23

jhonyy9 (jhonyy9):

but before all y not can being -5 and 5

OpenStudy (anonymous):

actually the equation is...\[9 \over y+5\]

OpenStudy (anonymous):

\[6 \over y-5\]

OpenStudy (anonymous):

\[=6 \over y^2-25\]

jhonyy9 (jhonyy9):

yes but you see that the denominator common will be y2 -25=(y-5)(y+5)

jhonyy9 (jhonyy9):

and hence the first fraction you need multiply by y-5 and second fraction by y+5

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

but the answer you have does not match up with any of the multiple choice answers

jhonyy9 (jhonyy9):

make the calcule newly

jhonyy9 (jhonyy9):

will be 9(y-5)-6(y+5)=6

jhonyy9 (jhonyy9):

so 9y-45-6y-30=6 3y-75=6 3y=81 y=27 ok i see the was what i have doit wrong sorry

jhonyy9 (jhonyy9):

now is right ?

OpenStudy (anonymous):

yes! thank you!

jhonyy9 (jhonyy9):

but you can prove it before accept like right answer

jhonyy9 (jhonyy9):

and please push the good answer button

jhonyy9 (jhonyy9):

thank you and good luck